simplify first or move variables all to one side. please show steps to solving the equation
5/4x-3/4=5/6x+1/2 *please help me I cant seem to get the right solution thank you!

Answer by stanbon(75887)

(5/4)x-(3/4) = (5/6)x+(1/2)
----------
Multiply thru by 12 to get:
15x - 9 = 10x + 6
5x = 15
x = 3
------------------
Cheers,
Stan H.